a dress shop owner has 8 new dresses that she wants to display in the window. If the display window has 5 mannequins, in how many ways can she dress them up?
Answer by CubeyThePenguin(3113)

Order matters, so we are talking about permutations.

(8!)/((8 - 5)!) = 8 * 7 * 6 * 5 * 4 = 6,720 ways